OpenStudy (anonymous):

find the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half

OpenStudy (anonymous):

Well, point-slope form is (y-y1)=m (x-x1) where M=slope. All you would have to do is substitue the values in, then simplify. (y-(-3)) =1/2 (x-6) (y+3)=1/2 (X-6)

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (anonymous):

No problem (: